# A Taxonomy of System-Level Attacks on Deep Learning Models in Autonomous Vehicles

### v1

@@ 4.1. Initial Paper Collection Through Scientific Databases
We employed Findpapers (grosman2020findpapers, ) to collect our initial set of papers. This tool allows us to design a customized search query, focusing on paper titles and abstracts, and supporting multiple databases including IEEE Explore [^2] https://ieeexplore.ieee.org/Xplore/home.jsp, Scopus [^3] https://www.elsevier.com/products/scopus, and ACM Digital Library [^4] https://dl.acm.org/. To ensure both comprehensive coverage and relevance, we iteratively refined our query based on two key criteria: the number of papers collected and the inclusion of three manually selected pivotal papers. Our pivotal papers contain two survey papers (biggio2018wild, ; cina2023wild, ) and DeepBillboard (zhou2020deepbillboard, ), which deal with adversarial attacks on autonomous vehicles [^5] These papers are excluded from the final taxonomy since they do not satisfy criteria introduced in later phases of the process (e.g., no survey; system-level failure propagation).. Through this iterative process, we balanced the query output, encompassing key papers while maintaining a broad search scope.
Given our focus on system-level attacks on autonomous vehicles, we designed a search query that requires the papers’ metadata (abstract, title, and keywords) to contain three key terms: ‘autonomous’, ‘vehicle’, and ‘attack’. Considering that not all papers may use these terms explicitly, we expanded our query to include synonyms. For ‘autonomous’, we include variations such as ‘automated’, ‘driverless’, ‘self-driving’, and ‘unmanned’, connected with the OR operator to account for any of these terms appearing in the metadata. Similarly, for ‘vehicle*’, we incorporate terms like ‘car*’, ‘drone*’, and ‘robot*’, also connected with OR. Note that the use of asterisks ensures coverage of plural forms. Lastly, for ‘attack*’, we include alternatives like ‘threat*’, ‘exploit*’, and ‘adversarial’, again joined by OR.
@@
([vehicle*] OR [car*] OR [drone*] OR [robot*])
@@
([attack*] OR [threat*] OR [exploit*] OR [adversarial]))
Executing this query across three databases from January 2017 to January 2024, we retrieved a total of 8,831 papers. Note that terms like ‘machine learning’. ’deep learning’ and ‘system-level’ are not explicitly mentioned in this query as they might exclude too many relevant papers. However, the subsequent manual filtering process ensured that only papers in the scope of our study (i.e., system level attacks against deep learning models) are taken into consideration. Moreover, the term ‘robots’ is included to incorporate dynamic and mobile autonomous robots, such as delivery robots.
Next, our task was to refine this extensive pool of papers by filtering out those considered irrelevant or lacking in quality. Initially, we removed duplicate papers. Subsequently, we evaluated the publication venues based on established academic standards, such as the presence of a rigorous peer-review process and their focus on topics relevant to security. Papers from venues that did not meet these criteria were excluded. Employing a venue frequency filter, we reduced the number of papers from 8,831 to 1,125. This filter was designed by ordering journals and conferences separately by frequency of occurrence in the pool and then selecting the top 50 most prevalent venues. Out of them, we eliminated venues lacking security-related keywords in their titles. For example, communication-related journals and conferences, and multimedia journals and conferences are excluded. Recognizing that this approach may exclude reputable software engineering/security-focused venues, we created a white list comprising prominent venues, ensuring their inclusion after the filtering process. This venue white-list is available online at https://zenodo.org/records/12806551.
@@ 4.2. Snowballing
The snowballing (wohlin2014guidelines, ) phase is aimed at expanding the pool of relevant papers by following their connections with other relevant papers. We employed both backward and forward snowballing techniques on the $3030$ papers obtained from the previous phase.
@@ 4.3. Creation of Taxonomy Categories
The process of category definition for the taxonomization of the papers was bootstrapped with the categories already available from the literature of attacks on DL models and autonomous vehicles. Initially, we adopted ten categories suggested by Biggio et al. (biggio2018wild, ) and Cinà et al. (cina2023wild, ). Subsequently, during taxonomy construction, we widened the scope to include concepts explicitly associated with system-level attacks and failures. This involved factors such as the system targeted by the attack and the specific system failures encountered during the attacks. In total, we identified 12 top-level categories, each listed in Table 1 with the corresponding explanations. A detailed description of these categories,a s weel as the sub-categories, is carried out below, in Section 5.3.
Table 1. Taxonomy Categories of System-level Attacks
@@
Application Domain The domain or type of vehicle in which the attacker is attempting to cause a system-level failure Cars, Drones, Robots
DL Model Under Attack The specific deep-learning model targeted by the attack Yolov5, ResNet-34, Dave2
System Under Attack The specific autonomous system targeted by the attack Baidu Apollo in LGSVL, Driving agent of CARLA
@@
In the phase of paper categorization and taxonomy refinement, we read each paper thoroughly. Specifically, each paper was read and categorized by at least two assessors. Weekly consensus meetings were held to resolve and negotiate any differences in the categorization. We excluded papers that do not specifically target the DL component of the AV under attack. For instance, papers that attack the hardware sensors and cause them to malfunction, resulting in incorrect point cloud generation and subsequent failure to recognize targets, are discarded as they do not directly involve any DL model (cao2023you, ). Additionally, papers that only evaluate their attack on a dataset without executing the system, or collect their own dataset using cameras mounted on non-autonomous, manually driven vehicles, are excluded (zhou2020deepbillboard, ). Papers that do not report any system-level failures, despite executing their attacks against autonomous vehicles (fu2022ad, ; muller2022physical, ), are also discarded. Although these papers may report mispredictions of the vehicle’s DL model, they do not provide information on the system-level vehicle’s behavior once the attack has caused a model-level misprediction. As a result of this strict refinement process, the number of relevant papers was reduced from $5050$ to $1919$.
@@ 5.1. Summary of papers
A summary of the 19 papers that met our criteria is provided here to give an overview before delving into the detailed classification of each paper. The papers are organized in chronological order, starting with the most recent to the oldest.
@@ 5.2. Mapping the papers
For the taxonomy mapping, we employed a rigorous process: each paper was independently reviewed and categorized by at least two assessors. To ensure consistency and reliability, the assessors regularly met to discuss and resolve disagreements in their mappings. The results of this process are presented in the Appendix, in Tables 2, 3, and 4, where each column represents a distinct category, and each row corresponds to an individual paper and its associated categorization.
@@ 5.3. Taxonomy tree
    Application Domain: Domain or vehicle type that is the target of the attack, such as “Cars” or “Drones”.
@@
    DL Model Under Attack: The DL model under attack, which is responsible for the system-level failure, such as YOLOv5, ResNet-34, or DAVE-2. The “DL Model Under Attack” category has four subcategories. The first, “Object Detection & Tracking”, includes DL models (e.g., YOLO) that handle object tracking or detection. The “Steering Wheel Angle Prediction” covers models that control the car’s steering angle. “Road Line Detection” helps the system detect and follow road lines, while “Traffic Control” assists in navigating complex traffic situations.
@@
    System Under Attack: The autonomous system under attack, such as a driving agent like Baidu Apollo in the LGSVL simulator or the driving agent in the CARLA simulator. “System Under Attack” has two subcategories, “Simulation” and “Real AVs”, respectively indicating that the system under attack is simulated or is a real-world autonomous vehicle. “Baidu Apollo” (apollo, ) and “CARLA’s Agent” (Dosovitskiy17, ) represent examples of autonomous systems implemented in the LGSVL (rong2020lgsvl, ) and CARLA simulators — the most popular simulation environments in our list of papers. Other less popular simulators include BeamNG (beamng_tech, ) and Udacity (udacity, ). “Tesla” is instead an example referred to by papers that used a real Tesla vehicle to test their attacks.
@@
    • “System-level White-Box”: The attacker has unrestricted access to the entire system, including the source code of the system, DL models, simulator (if any) and datasets.
@@
    Attack/Error Specificity: Some attacks are highly specific, targeting only one particular object or aspect (e.g., only traffic signs are attacked), which makes them “Attack-Specific”. On the other hand, the “Attack-Generic” subcategory represents a more general type of attack, which can be applied to a wider range of targets (e.g., any object on the side of the road). Regarding the model mis-predictions resulting from the attack, “Error-Specific” attacks aim to induce a particular output or trigger unique misbehavior in the model (e.g., confusing a 60 speed limit with a 90), while “Error-Generic” attacks allow any kind of misbehavior of the model (e.g., mis-classifying a traffic sign in any possible way).
@@ 6.1.1. RQ1 [Attack features]
In Figure 3(a), we observe that only two papers focus on the application domain of autonomous drones, compared to 17 papers on autonomous cars. As autonomous drones are gaining traction in several sectors, the unbalanced distribution of existing papers points to a whole research area that could be investigated more deeply in the future.
@@
Figure 3(b) shows that most attacks target the object detection and tracking models of AVs, indicating that this component is the most vulnerable part of an AV system. Steering wheel angle prediction was also a notable target for attacks, as it is influenced by numerous factors on the road that can be the target of an attack (von2023deepmaneuver, ; piazzesi2021attack, ; sato2021dirty, ; pavlitskaya2020feasibility, ; liu2017trojaning, ).
@@
Figure 3(c) clearly shows that the majority of papers focused on deploying attacks in a simulation environment. This trend may be due to the inherent dangers of the attacks, or legal constraints in various countries, making simulation environments more feasible for conducting such research.
@@
Figure 3(d) illustrates the approach chosen by attackers to deploy their attacks. Evasion attacks are prevalent, likely because poisoning attacks require access to the model’s dataset before training. This scenario is hard to achieve in the real world, making evasion attacks more realistic and feasible.
@@
In Figure 3(e), the model-level knowledge is predominantly considered to be white-box. While this may not be very practical in real-world scenarios, it is often necessary for attackers to make their attacks effective. In this respect, the four papers (piazzesi2021attack, ; jing2021too, ; boloor2020attacking, ; nassi2020phantom, ) that utilized a black-box knowledge approach would represent more feasible options for real-world attacks.
@@
Figure 3(f) addresses the system-level knowledge required for attacks. It is noteworthy that almost every attack does not require in-depth knowledge or access to the source code of the system. On one hand, this might be due to the practical difficulties that prevent the acquisition of white-box or grey-box knowledge of complex, possibly proprietary, systems. On the other hand, while at the model level, white-box knowledge is often a prerequisite that makes the attack otherwise impossible, at the system level plain knowledge of the environment elements and system input is often enough to introduce manipulations that cause a model mis-prediction and possibly a system failure.
@@
In line with Figure 3(b), Figure 3(g) also shows that the outcome of attacks on the DL component is predominantly the misdetection of targets or objects the AV is trying to recognize. These objects can include traffic signs, cars, pedestrians, etc. This highlights a significant vulnerability in the object detection and recognition capabilities of autonomous vehicles.
@@
Figure 3(h) suggests that most crashes involving AVs were with objects such as obstacles on the road. The figure indicates that only one paper reported crashes involving pedestrians. This could be related to the difficulty in effectively targeting pedestrians as part of an attack.
{mdframed}[backgroundcolor=cyan!15] RQ1 [Attack Features]: The distribution of papers across taxonomy categories is quite unbalanced. Most papers are focused on attacks against the object detection/tracking and the steering angle prediction components of simulated cars aiming for mispredictions (evasion). Most papers require white-box knowledge of the model, but black-box knowledge of the system. At the model level, they mostly aim at object misdetection, while at the system level, they mostly aim at crashing other vehicles or existing objects.

### v2

@@ 5.1. Initial Paper Collection Through Scientific Databases
We employed Findpapers (grosman2020findpapers, ) to collect our initial set of papers. This tool allows us to design a customized search query, focusing on paper titles and abstracts, and supporting multiple databases including IEEE Explore [^4] https://ieeexplore.ieee.org/Xplore/home.jsp, Scopus [^5] https://www.elsevier.com/products/scopus, and ACM Digital Library [^6] https://dl.acm.org/. To ensure both comprehensive coverage and relevance, we iteratively refined our query based on two key criteria: the number of papers collected and the inclusion of three manually selected pivotal papers. Our pivotal papers contain two survey papers (biggio2018wild, ; cina2023wild, ) and DeepBillboard (zhou2020deepbillboard, ), which deal with adversarial attacks on AVs [^7] These papers are excluded from the final taxonomy since they do not satisfy criteria introduced in later phases of the process (e.g., no survey; system-level failure propagation).. Through this iterative process, we balanced the query output, encompassing key papers while maintaining a broad search scope.
Given our focus on system-level attacks on AVs, we designed a search query that requires the papers’ metadata (abstract, title, and keywords) to contain three key terms: ‘autonomous’, ‘vehicle’, and ‘attack’. Considering that not all papers may use these terms explicitly, we expanded our query to include synonyms. For ‘autonomous’, we include variations such as ‘automated’, ‘driverless’, ‘self-driving’, and ‘unmanned’, connected with the OR operator to account for any of these terms appearing in the metadata. Similarly, for ‘vehicle*’, we incorporate terms like ‘car*’, ‘drone*’, ‘truck*’, ‘marine’, etc, also connected with OR. Note that the use of asterisks ensures coverage of plural forms.
@@
([vehicle*] OR [car*] OR [drone*] OR [truck*] OR [tractor*] OR [submarine*] OR [boat*] OR [maritime] OR [robotaxi*] OR [marine] OR [space] OR [spacecraft] OR [agriculture] OR [agricultural] OR [delivery] OR [construction])
@@
([attack*] OR [adversarial]))
Executing this query across three databases from January 2017 to January 2024, we retrieved a total of 9,748 papers. Note that terms like ‘machine learning’, ‘deep learning’ and ‘system-level’ are not explicitly mentioned in this query as they might exclude too many relevant papers. However, the subsequent manual filtering process ensured that only papers in the scope of our study (i.e., system level attacks against deep learning models) were taken into consideration. Moreover, the term ‘robots’ is included to incorporate dynamic and mobile autonomous robots, such as delivery robots.
Next, our task was to refine this extensive pool of papers by filtering out those considered irrelevant or lacking in quality. Initially, we removed duplicate papers. Subsequently, we evaluated the publication venues based on established academic standards, such as the presence of a rigorous peer-review process and their focus on topics relevant to security. Papers from venues that did not meet these criteria were excluded. Employing a venue frequency filter, we reduced the number of papers from 9,748 to 1,125. This filter was designed by ordering journals and conferences separately by frequency of occurrence in the pool and then selecting the top 50 most prevalent venues. Out of them, we eliminated venues lacking security-related keywords in their titles. For example, communication-related journals and conferences, and multimedia journals and conferences are excluded. Recognizing that this approach may exclude reputable software engineering/security-focused venues, we created a white list comprising prominent venues, ensuring their inclusion after the filtering process. This venue white-list is available online at https://zenodo.org/records/16407709.
@@ 5.2. Snowballing
The snowballing (wohlin2014guidelines, ) phase is aimed at expanding the pool of relevant papers by following their connections with other relevant papers. We employed both backward and forward snowballing techniques on the $30$ papers obtained from the previous phase.
@@ 5.3. Creation of Taxonomy Categories
The process of category definition for the taxonomization of the papers was bootstrapped with the categories already available from the literature of attacks on DL models and AVs. Initially, we adopted ten categories suggested by Biggio et al. (biggio2018wild, ) and Cinà et al. (cina2023wild, ). Subsequently, during taxonomy construction, we widened the scope to include concepts explicitly associated with system-level attacks and failures. This involved factors such as the system targeted by the attack and the specific system failures encountered during the attacks. In total, we identified 12 top-level categories, each listed in Table 2 with the corresponding explanations. A detailed description of these categories, as well as the sub-categories, is carried out below, in Section 6.3.
Table 2. Examples of Possible Values of Each Taxonomy Category of System-level Attacks
@@
Application Domain The domain or type of vehicle in which the attacker is attempting to cause a system-level failure Cars, Drones, Trucks, Ships
DL Model Under Attack The specific deep-learning model and module, targeted by the attack Yolov5, ResNet-34, Dave2
System Under Attack The specific autonomous system targeted by the attack Baidu Apollo in LGSVL, CARLA’s agents (MSF or SSF)
@@
In the phase of paper categorization and taxonomy refinement, we read each paper thoroughly. Specifically, each paper was read and categorized by at least two assessors. Weekly consensus meetings were held to resolve and negotiate any differences in the categorization. We excluded papers that do not specifically target the DL component of the AV under attack. For instance, papers that attack the hardware sensors and cause them to malfunction, resulting in incorrect point cloud generation and subsequent failure to recognize targets, are discarded as they do not directly involve any DL model (cao2023you, ). Additionally, papers that only evaluate their attack on a dataset without executing the system, or collect their own dataset using cameras mounted on non-autonomous, manually driven vehicles, are excluded (zhou2020deepbillboard, ). This criterion led to the exclusion of many papers, including the only two studies that specifically targeted Maritime Autonomous Surface Ships (MASS) (lee2023eval, ; lee2023vuln, ). Papers that do not report any system-level failures, despite executing their attacks against AVs (fu2022ad, ; muller2022physical, ; huang2024spot, ), are also discarded. Although these papers may report mispredictions of the vehicle’s DL model, they do not provide information on the system-level vehicle’s behavior once the attack has caused a model-level misprediction. As a result of this strict refinement process, the number of relevant papers was reduced from $50$ to $21$. Other avenues for understanding attacks on AVs include research that targets network-based functionalities or attacks on DL models without system-level evaluation, which were not our goal in this taxonomy.
@@ 6.1. Summary of papers
A summary of the 21 papers that met our criteria is provided here to give an overview before delving into the detailed classification of each paper. The papers are organized in chronological order, starting with the most recent to the oldest.
@@
“RPAU: Fooling the Eyes of UAVs via Physical Adversarial Patches”: Liu et al. (liu2024rpau, ) placed adversarial patches in the environment, including on moving vehicles, to disrupt drone perception. These patches caused the drones to either miss real objects and crash into them, change direction and veer off course, or detect fake objects and freeze in response.
“Adversarial Attacks on Adaptive Cruise Control Systems”: Guo et al. (guo2023adv, ) used leading vehicles with adversarial patches placed on their rear, causing the target vehicle to either fail to detect the vehicle ahead or perceive it as being farther away – potentially leading to a collision. To make the attack stealthy, they disguised the patch as a fake advertisement displayed on the back of a truck.
@@ 6.2. Mapping the papers
For the taxonomy mapping, we employed a rigorous process: each paper was independently reviewed and categorized by at least two assessors. To ensure consistency and reliability, the assessors regularly met to discuss and resolve disagreements in their mappings. The results of this process are presented in the Appendix, in Tables 3, 4, and 5, where each column represents a distinct category, and each row corresponds to an individual paper and its associated categorization.
@@ 6.3. Taxonomy tree
    Application Domain: Domain or vehicle type that is the target of the attack, such as “Cars” and “Drones. If we compare the application domains covered by the papers included in the taxonomy with the AV types reported by the LLM in Table 1, we can notice that most AV surfaces, including sea, underwater and space, are uncovered. On the other hand, all these uncovered AV types represent critical domains where security testing would be quite important. We conclude that research is missing on the system level security of AVs operating on surfaces other than land and air.
@@
    DL Model Under Attack: The DL model under attack, which is responsible for the system-level failure, such as YOLOv5, ResNet-34, or DAVE-2. The “DL Model Under Attack” category has four subcategories. The first, “Object Detection & Tracking”, includes DL models (e.g., YOLO) that handle object tracking or detection. The “Steering Wheel Angle Prediction” covers models that control the car’s steering angle. “Road Line Detection” helps the system detect and follow road lines, while “Traffic Control” assists in navigating complex traffic situations. All of these DL models contribute to the autonomous capabilities of vehicles and are part of a modular AV architecture, such as perception, planning, and end-to-end driving. If we compare the models under attack covered by the papers included in the taxonomy with the AV models and modules reported by the LLM in Table 1, we can notice that most types of modules are covered. However, modules for localization and mapping have not been subjected to system-level security testing in any of the included papers. This might be due to the difficulty associated with faithful simulation of such modules. Another missing module type is associated with the low level “control” functionality, often implemented via reinforcement learning. This indicates a general preference of security testing papers toward supervised learning, as opposed to reinforcement learning. A similar preference seems to hold for model-level adversarial attacks, which are typically crafted against neural networks trained in a supervised setting.
@@
    System Under Attack: The autonomous system under attack, such as a driving agent like Baidu Apollo in the LGSVL simulator or the driving agent in the CARLA simulator. “System Under Attack” has two subcategories, “Simulation” and “Real AVs”, respectively indicating that the system under attack is simulated or is a real-world autonomous vehicle. “Baidu Apollo” (apollo, ) and “CARLA’s Agent” (Dosovitskiy17, ) represent examples of autonomous systems implemented in the LGSVL (rong2020lgsvl, ) and CARLA simulators — the most popular simulation environments in our list of papers. Other less popular simulators include BeamNG (beamng_tech, ) and Udacity (udacity, ). “Tesla” is instead an example referred to by papers that used a real Tesla vehicle to test their attacks. As systems in AVs may use either single or multi-sensor fusion to make autonomous decisions, in this category, we distinguish between them by using MSF for multi-sensor fusion systems and SSF for single-sensor fusion systems. If we compare the systems under attack covered by the papers included in the taxonomy with the simulation environments reported by the LLM in Table 1, we can notice that simulators for several AV surfaces, including sea, underwater and space, are not used in the included papers, or are used for another surface (e.g., Gazebo can be used to simulate air, sea, underwater and space, but in the included papers it was used only to simulate AVs operating in the air – specifically, drones). Similarly to the remark made above for the application domain, the missing simulation environments point to the need for further research to bring system level attacks to simulators operating in the sea, underwater or in the space.
@@
    • “System-level White-Box”: The attacker has unrestricted access to the entire system, including the source code of the system, the hardware specs, DL models, simulator (if any) and datasets.
@@
    Attack/Error Specificity: This category includes Attack-Specific or Attack-Generic, as well as Error-Specific or Error-Generic. Some attacks are highly specific, targeting only one particular object or aspect (e.g., only traffic signs are attacked), which makes them “Attack-Specific”. On the other hand, the “Attack-Generic” subcategory represents a more general type of attack, which can be applied to a wider range of targets (e.g., any object on the side of the road). Regarding the model mis-predictions resulting from the attack, “Error-Specific” attacks aim to induce a particular output or trigger unique misbehavior in the model (e.g., confusing a 60 speed limit with a 90), while “Error-Generic” attacks allow any kind of misbehavior of the model (e.g., mis-classifying a traffic sign in any possible way).
@@ 7.1.1. RQ1 [Attack features]
Figure 3(a) illustrates that the perception module is the most frequently targeted in attacks (ma2024slowtrack, ; cao2019adversarial, ; yoon2023learning, ; hanfeld2023kidnapping, ; wang2023does, ; zhang2023data, ; yan2022rolling, ; cao2021invisible, ; jing2021too, ; yang2021robust, ; jha2020ml, ; nassi2020phantom, ; guo2023adv, ). End-to-end models follow as the second most common target (liu2024rpau, ; von2023deepmaneuver, ; piazzesi2021attack, ; sato2021dirty, ; boloor2020attacking, ; pavlitskaya2020feasibility, ; liu2017trojaning, ). However, only a single attack was recorded on the prediction and planning module, while no attacks were reported on the DL components of the localization or control modules. Since the perception module is the first to process sensor data in an AV, any compromise at this stage can lead downstream modules to make incorrect decisions, ultimately resulting in faulty control outputs. End-to-end models are also highly vulnerable, as they bypass preprocessing and act directly on raw sensor inputs, increasing their susceptibility to adversarial manipulation.
Figure 3(b) shows that attacks are fairly evenly distributed between MSF and SSF systems. Although MSF systems are often considered more secure due to their integration of multiple sensor types (lou2022testing, ), this study reveals that they remain vulnerable to well-crafted adversarial attacks.
@@
In Figure 3(c), we observe that only three papers focus on the application domain of autonomous drones, compared to 18 papers on autonomous cars. As autonomous drones continue to gain traction across various sectors, we expect such proportion to change in the future. The notable lack of research on system-level adversarial attacks in other application domains, such as autonomous trucks, buses, ships, spacecraft, and more, highlights a clear gap in the current literature, pointing to a whole research area that could be investigated more deeply in the future.
@@
Figure 3(d) shows that most attacks target the object detection and tracking models of AVs, indicating that this component is the most vulnerable part of an AV system. Steering wheel angle prediction was also a notable target for attacks, as it is influenced by numerous factors on the road that can be the target of an attack (von2023deepmaneuver, ; piazzesi2021attack, ; sato2021dirty, ; pavlitskaya2020feasibility, ; liu2017trojaning, ).
@@
Figure 3(e) clearly shows that the majority of papers focused on deploying attacks in a simulation environment. This trend may be due to the inherent dangers of the attacks, or legal constraints in various countries, making simulation environments more feasible for conducting such research.
@@
Figure 3(f) illustrates the approach chosen by attackers to deploy their attacks. Evasion attacks are prevalent, likely because poisoning attacks require access to the model’s dataset before training. This scenario is hard to achieve in the real world, making evasion attacks more realistic and feasible.
@@
In Figure 3(g), the model-level knowledge is predominantly considered to be white-box. While this may not be very practical in real-world scenarios, it is often necessary for attackers to make their attacks effective. In this respect, the four papers (piazzesi2021attack, ; jing2021too, ; boloor2020attacking, ; nassi2020phantom, ) that utilized a black-box knowledge approach would represent more feasible options for real-world attacks.
@@
Figure 3(h) addresses the system-level knowledge required for attacks. It is noteworthy that almost every attack does not require in-depth knowledge or access to the source code of the system. On one hand, this might be due to the practical difficulties that prevent the acquisition of white-box or grey-box knowledge of complex, possibly proprietary, systems. On the other hand, while at the model level, white-box knowledge is often a prerequisite that makes the attack otherwise impossible, at the system level, plain knowledge of the environment elements and system input is often enough to introduce manipulations that cause a model mis-prediction and possibly a system failure.
@@
In line with Figure 3(d), Figure 3(i) also shows that the outcome of attacks on the DL component is predominantly the misdetection of targets or objects the AV is trying to recognize. These objects can include traffic signs, cars, pedestrians, etc. This highlights a significant vulnerability in the object detection and recognition capabilities of autonomous vehicles.
@@
Figure 3(j) suggests that most crashes involving AVs were with objects such as obstacles on the road. The figure indicates that only one paper reported crashes involving pedestrians. This could be related to the difficulty in effectively targeting pedestrians as part of an attack.
RQ1 [Attack Features]: The distribution of papers across the taxonomy categories is notably unbalanced. System-level attacks on AVs are primarily focused on cars and, to a lesser extent, drones. Other types of AVs, such as tractors, spacecraft, and maritime surface ships, are largely absent from the current research landscape. Most attacks target the perception module of AVs, with significantly less attention given to other critical components like planning, localization, and control. While one might assume that AVs employing MSF are more resilient to adversarial attacks, this taxonomy reveals that MSF systems are still vulnerable. The majority of studies focus on simulated car environments, targeting components such as object detection/tracking and steering angle prediction, with the primary goal being mispredictions (evasion). These attacks are often conducted with white-box access to the model but only black-box knowledge of the system. At the model level, attacks are typically designed to cause object misdetection, whereas at the system level, the aim is often to cause the AV to crash into other vehicles or obstacles.